OpenStudy (ujjwal):

A parallel beam of sodium light is incident normally on a diffraction grating. The angle between the two first order spectra on either side of normal is 27 degree 42 minute . Assuming that the wavelength of light is 5.893* 10^-7 m. Find the number of rulings per mm on the grating.

OpenStudy (experimentx):



OpenStudy (ujjwal):

And how does that figure give the solution?

OpenStudy (experimentx):

we have relation, \( d \sin\theta = n \lamda \)

OpenStudy (experimentx):

for first order, n=1, since \( \theta\) is given, we can calculate d, but i don't understand why two results are given.

OpenStudy (ujjwal):

and what do we do with d? well the answer is 406 per mm.

OpenStudy (ujjwal):

d can't have that greater value..

OpenStudy (experimentx):

procedure seems all right http://www.physics.smu.edu/~scalise/emmanual/diffraction/lab.html

OpenStudy (ujjwal):

there was a calculation error earlier. And i think 27 degree and 42 min should be divided by 2 to get actual value of theta.. since the angle given in question is the angle between TWO first order spectra.. the calculations then made by me gives a value of 2.46*10^-6 for d. suppose i am correct. what next?

OpenStudy (experimentx):

d is the distance between two adjacent grating lines, change d into mm and divide 1 by this value.

OpenStudy (experimentx):

decreasing value of angel gives crazy data.

OpenStudy (ujjwal):

thanks! i got the answer!!

OpenStudy (experimentx):

how??

OpenStudy (ujjwal):

the same formula. divided the angle given in question by 2 so, theta= 13 degree 51 min, lambda is given, n=1.. that makes d=2.46*10^-6 m. converted this value into mm.. and the reciprocal of then obtained value was the answer.. you made calculations mistake earlier. however, formula was useful. thanks buddy!

OpenStudy (experimentx):

haha .. sure you are welcome.